# EEG-ImageNet: An Electroencephalogram Dataset and Benchmarks with Image Visual Stimuli of Multi-Granularity Labels

### v1

@@ 3.5 Dataset Description
The EEG-ImageNet dataset contains a total of 63,850 EEG-image pairs from 16 participants. Each EEG data sample has a size of (n_channels, ), where n_channels is the number of EEG electrodes, which is 62 in our dataset; is the sampling frequency of the device, which is 1000 Hz in our dataset; and T is the time window size, which in our dataset is the duration of the image stimulus presentation, i.e., 0.5 seconds. Due to ImageNet’s copyright restrictions, our dataset only provides the file index of each image in ImageNet and the wnid of its category corresponding to each EEG segment. Additional information about the dataset is shown in Appendix A.1.
4 Benchmarks
In this section, we detail the benchmarks of our study shown in Figure 1 by outlining the preprocessing steps, feature extraction methods, task definitions, and models used.
@@ 4.1 Preprocessing
We perform a series of preprocessing steps for the raw EEG data we collect to eliminate noise and artifacts and improve signal quality. The preprocessing pipeline includes the following stages: First, re-referencing: Re-referencing is done using the offline linked mastoids method, which uses the average of the M1 and M2 mastoid electrodes as the new reference point [39]. This minimizes potential biases and improves the signal-to-noise ratio. Then, filtering: Filtering is performed using a 0.5 Hz to 80 Hz band-pass filter to remove low-frequency drifts (<0.5 Hz) and high-frequency noise (>80 Hz). Additionally, 50 Hz environmental noise is eliminated. Finally, artifact removal: Artifact removal eliminates abnormal amplitude signals and artifacts caused by blinks or head movements.
@@ 4.2 Feature Extraction
In our benchmarks, for models requiring time-domain signals as direct input, we extract the 40ms-440ms segment of each EEG signal as the feature input. This approach helps to minimize the influence of preceding and subsequent image stimuli on the current stimulus. For models requiring frequency-domain features as input, we extract the differential entropy (DE) of the extracted time-domain signals as features, as this characteristic effectively captures the complexity and variability of brain activity in the frequency domain [7]. According to the general division in neuroscience, the frequency bands are categorized as delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-13 Hz), beta (13-30 Hz), and gamma (30-80 Hz). We use the Welch method with a sliding window to estimate the power spectral density in each frequency band. Then, we normalize the data and calculate the differential entropy (DE) using the formula, . Consequently, for each segment of EEG signals, we obtain the differential entropy (DE) for each electrode and each frequency band.
@@ 4.3 Task Definition
In our benchmarks, we test our dataset on two tasks: object classification and image reconstruction. The object classification task aims to classify the category of the corresponding image stimulus the participant is exposed to with their EEG signals. We evaluate the models using classification accuracy. In the image reconstruction task, given a specific EEG segment, the goal is to reconstruct the image stimulus the participant is exposed to. We evaluate the generated results using two-way identification [30] under different visual neural networks (refer to Appendix A.3 for details.).
On our multi-granularity labeled image dataset, we test various tasks with different levels of granularity. Additionally, due to the inherent significant inter-individual differences in EEG signals, all models in our benchmarks are trained exclusively in an intra-subject experimental setup. Furthermore, to mitigate the significant temporal effects [19] observed in the EEG-Classification dataset, all our experimental setups strictly adhere to a dataset split where the first 30 images of each category are used as the training set, and the last 20 images are used as the test set. We strongly recommend that all researchers using the EEG-ImageNet dataset adopt a similar dataset split methodology. We also explain the impact of this split in Appendix A.5.
@@ 4.4 Models
In the object classification task, we employ simple machine-learning classification models such as ridge regression, KNN, random forest, and SVM. Additionally, we implement deep learning models including MLP, EEGNet [17], and RGNN [44]. MLP consists of two hidden layers, while EEGNet and RGNN are implemented using their original architectures. We use cross-entropy as the loss function. We train the models for each participant for 2000-3000 epochs on an NVIDIA 4090 GPU.
In the image reconstruction task, we use a frozen Stable Diffusion 1.4 [29] as the backbone and train a two-layer MLP as an encoder to generate the prompt embeddings for input. For our reconstructions, we use 50 denoising timesteps with PNDM noise scheduling to generate 512x512 images. We use MSE as the loss function to align the encoder’s output with the CLIP (ViT-L/14) [26] embeddings of the captions of real images obtained through BLIP [18]. The model for each participant is trained for 2000 epochs with on an NVIDIA 4090 GPU.
For model and training details, please refer to the open-source code and Appendix A.3.
5 Experiments
In this chapter, we present the experimental results of our study, focusing on two benchmark tasks: object classification and image reconstruction.
5.1 Object Classification
In the object classification task, we conduct experiments at three levels of granularity: all, coarse, and fine. The “all” category represents the 80-class classification accuracy, the “coarse” category represents the 40-class classification accuracy, and the “fine” category represents the average accuracy of five 8-class classification tasks. The performance of each model is detailed in Table 2, which shows the average results of all participants.
| Model | Acc (all) | Acc (coarse) | Acc (fine) | |
| Classic model | Ridge | 0.2859† | 0.3944† | 0.5833† |
| KNN | 0.3037† | 0.4012† | 0.6954† | |
| RandomForest | 0.3489† | 0.4535† | 0.7288† | |
| SVM | 0.3919 | 0.5057† | 0.7784† | |
| Deep model | MLP | 0.4037 | 0.5339 | 0.8163 |
| EEGNet* | 0.2604† | 0.3030† | 0.3645† | |
| RGNN | 0.4050 | 0.4703† | 0.7057† | |
The table shows that RGNN achieves the highest accuracy in the “all” task among all models, with an 80-class classification accuracy reaching 40.50%. MLP slightly lags behind RGNN in the “all” task accuracy, but it significantly outperforms in the “coarse” and “fine” tasks, with a 40-class accuracy of 53.39% and an average 8-class accuracy of 81.63%. Among the classical machine learning models, SVM exhibits the highest classification performance, only slightly trailing RGNN and MLP across the three different granularity scenarios. The lower accuracy of RGNN in the “fine” task might be due to the complex model structure, which may require more fine-tuning to better adapt to easier tasks. Meanwhile, the lower performance of time-domain features compared to frequency-domain features suggests that frequency-domain features might be more informative for the classification tasks in this study. Specific performance details for each participant in each task are provided in Appendix A.4. We find that the ranking of participants’ accuracy is relatively consistent across different models.
@@ 5.1 Object Classification
We compile the accuracy of each participant for the “all” task under the SVM, MLP, and RGNN models, as shown in Figure 2(a). We observe significant differences between participants, indicating high variability in individual responses. The best-performing participant achieves an accuracy of 60.88%. To better compare the differences between “fine” and “coarse” tasks in EEG-ImageNet, we modify the coarse-grained task. We randomly select 8 coarse-grained categories and use the RGNN model for training and testing. This process is repeated 5 times, and the average accuracy for each participant is calculated and plotted alongside their average “fine” task accuracy in Figure 2(b). We then perform linear regression on the data points, and the resulting function has a slope greater than 1, indicating that participants generally achieve better results on the “coarse” classification tasks. This finding is also consistent with intuition.
5.2 Image Reconstruction
Figure 3 shows some of the results from our image reconstruction pipeline of a single participant (S8). From the selected images, it can be seen that the reconstruction pipeline can effectively restore the category information of the image stimuli. However, restoring low-level details such as color, position, and shape is inaccurate. This may be because we align the EEG signals to the image captions obtained from BLIP. Due to the limited descriptive precision of BLIP, we cannot fully leverage the diffusion model’s generative capabilities.
Table 3 shows the average two-way identification (chance=50%) of the images generated by our reconstruction pipeline using different visual neural networks. Two-way identification refers to comparing the generated image with the original and one distractor to evaluate accuracy. The highest accuracy achieved by CLIP suggests that integrating vision transformers with extensive pre-training on diverse image-text data can significantly enhance model performance for complex tasks like image reconstruction from EEG signals. Additional results can be found in Appendix A.4.
| Method | Alex(2) | Alex(5) | Incep | CLIP(ViT-L/14) |
| Two-way Identification | 56.05% | 62.99% | 56.75% | 64.67% |
@@ 6 Discussion and Conclusion
In this section, we outline the limitations of our dataset and explore how it could guide future research efforts to advance machine learning and brain-computer interface design.
@@
Limitation. Firstly, while our dataset is more comprehensive than similar works, each participant’s data is still relatively limited. This necessitates the development of inter-subject models to overcome this limitation and enhance generalizability. Additionally, it is limited in representation, as participants were drawn from a convenience sample at our university. This results in an age distribution skewed towards college-aged individuals and a racial composition predominantly White and Asian, which limits the dataset’s generalizability. Future work should aim to include a more diverse and extensive participant pool. Secondly, while we employed methods such as reducing the segment length of each EEG recording and sequentially splitting the training and test sets to mitigate the temporal effect, we were unable to eliminate it completely. Future work should explore more sophisticated techniques to address this issue. Lastly, our benchmarks did not incorporate many of the latest deep-learning methods. We believe that recent advancements in deep learning could greatly benefit from our comprehensive dataset, potentially leading to significant breakthroughs in visual neuroscience.
@@
Insight for ML. The EEG-ImageNet dataset provides a comprehensive resource for developing models in visual recognition tasks, enabling the development of sophisticated deep-learning models capable of capturing intricate patterns within EEG data. Future research could leverage the dataset to enhance domain adaptation and transfer learning techniques, facilitating effective inter-subject task completion. Researchers might develop state-of-the-art models for the benchmarks we defined, or even create new tasks. By offering a diverse set of visual stimuli and supporting multi-level classification tasks, EEG-ImageNet could foster the creation of hierarchical models that mirror human cognitive processes and improve the generalization capabilities of machine learning algorithms.
@@
Insight for BCI. As hardware technology progresses, portable EEG devices are becoming increasingly feasible, offering new opportunities for real-time BCI applications. Researchers could use the dataset to develop robust BCI systems that accurately interpret user intent from EEG signals. The comprehensive size and diverse visual stimuli in EEG-ImageNet allow for the creation of adaptive BCI systems that learn and respond to individual user patterns. This paves the way for personalized neurotechnology solutions, particularly enhancing human-computer interaction for individuals with disabilities. Furthermore, addressing privacy protection and ethical concerns will be crucial as BCI technology advances, ensuring user data is securely handled and individual rights are respected.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] See Section 6.
- (c)
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Section 3.1.
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [TODO]
Did you include complete proofs of all theoretical results? [TODO]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 3.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 4.
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] The reported experimental results were averaged across multiple participants after being conducted for each participant individually.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] We use images from ImageNet as visual stimuli and cite relevant works.
Did you mention the license of the assets? [Yes] In the GitHub repository.
Did you include any new assets either in the supplemental material or as a URL? [Yes] In the GitHub repository.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] We registered to access ImageNet21k, and due to copyright protection, we only annotated the dataset with the image indices and the corresponding categories’ wnids.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section 3.1.
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes] See Section 3.1.
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Section 3.
@@ A.1 Additional Information about Dataset
| EEG-ImageNet | 80 | 4000 | 16 | 63850 | 15.88GB |
As shown in Listing 1, the EEG-ImageNet dataset storage format is provided. The dataset can be accessed through the cloud storage link available in our GitHub repository https://github.com/Promise-Z5Q2SQ/EEG-ImageNet-Dataset. Due to file size limitations on the cloud storage platform, we split the dataset into two parts: “EEG-ImageNet_1.pth” and “EEG-ImageNet_2.pth”, each containing data from 8 participants. Users can choose to use only one of the parts based on their specific needs or device limitations.

### v2

@@ 3.4 Dataset Description
The CrossPT-EEG dataset contains a total of 87,850 EEG-image pairs from 16 participants, with a total of 22 sessions (6 participants took part in two sessions). Each EEG data sample has a size of (, ), where is the number of EEG electrodes, which is 62 in our dataset; is the sampling frequency of the device, which is 1000 Hz in our dataset; and T is the time window size, which in our dataset is the duration of the image stimulus presentation, i.e., 500 ms. Due to ImageNet’s copyright restrictions, our dataset only provides the file index of each image in ImageNet and the wnid of its category corresponding to each EEG segment. Additional information about the dataset is shown in Appendix A.1.
4 Benchmarks Settings
In this section, we detail the benchmarks of our study by outlining the preprocessing steps, feature extraction methods, task definitions, and models used. The detailed code can be accessed openly through the url https://anonymous.4open.science/r/EEG-ImageNet-anonymous-74B0.
@@ 4.1 Preprocessing
We perform a series of preprocessing steps for the raw EEG data we collect to eliminate noise and artifacts and improve signal quality (ye2024relevance). The preprocessing pipeline includes the following steps: First, re-referencing is done using the offline linked mastoids method, which uses the average of the M1 and M2 mastoid electrodes as the new reference point (yao2019reference). Then, filtering is performed using a 0.5 Hz to 80 Hz band-pass filter to remove low-frequency drifts and high-frequency noise. Additionally, 50 Hz environmental noise is eliminated. Finally, artifact removal eliminates abnormal amplitude signals and artifacts caused by blinks or head movements.
@@ 4.2 Feature Extraction
In our benchmarks, we extract the 40ms-440ms segment of each EEG signal as the domain feature input. This approach helps to minimize the influence of preceding and subsequent image stimuli on the current stimulus. For models requiring frequency-domain features as input, we extract the differential entropy (DE) of the extracted time-domain signals as features, as this characteristic effectively captures the complexity and variability of brain activity in the frequency domain (duan2013differential). According to the general division in neuroscience, the frequency bands are categorized as delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-13 Hz), beta (13-30 Hz), and gamma (30-80 Hz). We use the Welch method with a sliding window to estimate the power spectral density in each frequency band. Then, we normalize the data and calculate the differential entropy (DE) using the formula, . Consequently, for each segment of EEG signals, we obtain the differential entropy (DE) for each electrode and each frequency band.
@@ 4.3 Task Definition
To demonstrate the effectiveness of CrossPT-EEG with its cross-time, two-stage design, we conduct EEG-based object classification tasks under different experimental settings. The object classification task is defined as predicting the category of the image stimulus that a participant is viewing, based solely on their EEG signals.
It is important to note that block-design (all images of the same category are presented together) experiments can introduce temporal markers into EEG data, which may artificially inflate classification performance when evaluated only within Stage 1. While such results are not representative of real-world generalization, they may still provide useful insights for future research on EEG temporal effects. For completeness, we include these Stage 1-only results in the Appendix A.4.
We conduct evaluations on EEG-based visual decoding with four tasks, Within-Time (WT), Cross-Time (CT), Cross-Participant (CP), and Pre-training (PT). As shown in Figure 2, WT using 30 images per category for training and 20 images per category for testing, both from Stage 2. As WT adopted a non-block-design paradigm in which images used for training and testing are temporally separated, the results are free from class-specific temporal bias. In the CT task, the training data come from Stage 1 of a given participant, while the testing data come from Stage 2 of the same participant. Thus, CT evaluates cross-time generalization within the same individual. In the CP task, the training data also come from Stage 1, but specifically from the 10 participants who did not participate in Stage 2. The testing data comes from Stage 2 of the remaining participant. Thus, CP evaluates generalization across participants as well as across sessions. Finally, we investigated cross-subject, cross-time pre-training experiments (PT) to investigate whether pre-training on other participants’ data could improve recognition performance in cross-time classification. This setting enables us to examine the benefits of leveraging additional subjects for pre-training in enhancing the robustness and generalization of visual decoding models. On our multi-granularity labeled image dataset, we test the above tasks with different levels of granularity.
In our experiments, we employed two-way identification as the primary evaluation metric (ozcelik2023natural). We chose this measure because it provides a clearer basis for cross-task comparisons. For completeness, we also report additional metrics, such as accuracy, in the Appendix A.4. In two-way identification, for each test sample, we randomly select k comparison samples from different categories. The model is then required to decide which of the two samples is more likely to belong to the ground-truth class. Each decision is scored as 1 if correct and 0 otherwise. Final performance is reported as the average over all pairwise comparisons, with a chance level of 0.5. In our benchmarks, we set k = 500 to ensure stable and robust evaluation.
@@ 4.4 Models
We employ simple machine-learning classification models such as ridge regression, KNN, random forest, and SVM. Additionally, we implement deep learning models including MLP, EEGNet (lawhern2018eegnet) and RGNN (zhong2020eeg). These models are the most commonly used and have demonstrated excellent performance in EEG-related research. We train the models on a single NVIDIA 4090 GPU. For model and training details, please refer to the code and Appendix A.3.
5 Experimental Results
| Model | WT | CT | CP | PT | |
| Classic model | Ridge | 0.861±0.022† | 0.509±0.011 | 0.501±0.007 | 0.508±0.010 |
| KNN | 0.892±0.026† | 0.517±0.019 | 0.491±0.006 | 0.521±0.017 | |
| RandomForest | 0.909±0.031† | 0.511±0.019† | 0.507±0.011 | 0.524±0.020† | |
| SVM | 0.935±0.039† | 0.536±0.018† | 0.506±0.015 | 0.544±0.017† | |
| Deep model | MLP | 0.938±0.024† | 0.552±0.012† | 0.524±0.016† | 0.570±0.020† |
| EEGNet* | 0.878±0.022† | 0.519±0.015† | 0.509±0.012 | 0.524±0.022† | |
| RGNN | 0.933±0.025† | 0.566±0.015† | 0.526±0.018† | 0.585±0.027† | |
| Model | WT | CT | CP | PT | |
| Classic model | Ridge | 0.886±0.023† | 0.549±0.017† | 0.495±0.007 | 0.543±0.012† |
| KNN | 0.909±0.025† | 0.554±0.017† | 0.505±0.010 | 0.551±0.015† | |
| RandomForest | 0.922±0.030† | 0.587±0.022† | 0.511±0.012 | 0.595±0.019† | |
| SVM | 0.949±0.035† | 0.592±0.021† | 0.509±0.019 | 0.606±0.022† | |
| Deep model | MLP | 0.954±0.026† | 0.617±0.019† | 0.563±0.026† | 0.636±0.027† |
| EEGNet* | 0.893±0.023† | 0.566±0.027† | 0.522±0.026† | 0.579±0.025† | |
| RGNN | 0.945±0.024† | 0.601±0.024† | 0.573±0.028† | 0.634±0.027† | |
The coarse-grained represents the 40-class classification accuracy; and the fine-grained represents the average accuracy of five 8-class classification tasks. Table 2 and Table 3 shows the average results of all participants in the coarse-grained and grained classification task, respectively. We observe that in the within-task (WT) task, both classic models and deep models achieve strong performance, indicating that WT is the relatively easiest among the four tasks. The best-performing model reaches a two-way identification close to 0.95, suggesting that EEG signals contain sufficient information for reliable decoding when training and testing are conducted within the same stage. However, in the cross-time (CT) task, performance drops substantially across all models. This degradation is particularly obvious for classic models, which primarily rely on learning fixed linear mappings. This result highlights the challenge posed by temporal variability in EEG signals. In the cross-participant (CP) task, the performance of classic models is further reduced, in some cases barely above chance, suggesting that these methods fail to generalize across individuals. By contrast, deep models consistently deliver better results across tasks. This indicates that graph-based neural networks are able to capture more stable and transferable patterns from EEG signals, making them promising candidates for addressing inter-individual differences. Meanwhile, the relatively poor performance of EEGNet may be attributed to the fact that the original version of EEGNet was not designed for semantically related tasks. Finally, in the pre-training (PT) task, deep models again show clear advantages, with RGNN and MLP achieving the highest performance across coarse-grained and fine-grained tasks. In the PT task, deep models outperform their CT counterparts across different levels of granularity, suggesting that leveraging pre-training with data from other participants can effectively improve cross-time recognition.
@@ 5 Experimental Results
To better compare the differences between “fine” and “coarse” tasks, we randomly select 8 coarse-grained categories and perform WT task because the result is the best on this task. This process is repeated 5 times, and the results for each participant is calculated and plotted alongside their average “fine” task results in Figure 3. We then perform linear regression on the data points, and the resulting function has a slope greater than 1, indicating that models generally achieve better results on the “coarse” classification tasks. This finding is also consistent with intuition.
We further aggregate results of different participants across different tasks for comparison, as shown in Figure 4. MLP and RGNN exhibit remarkably similar performance distributions across different tasks under both coarse and fine granularity settings. Both models follow the same trend ordering: . This indicates that within-time classification remains the easiest task, while cross-participant is consistently the most challenging. This trend aligns with prior works that have compared cross-subject and cross-session EEG decoding, which often report that the most difficult scenario is generalization across participants (liang2022multi; chen2021ms; apicella2024toward). Our results not only corroborate this conventional wisdom but also further demonstrate that pre-training on data from other participants can significantly boost object classification performance relative to direct cross-time generalization.
@@ 6 Discussion and Conclusion
In this paper, we introduced CrossPT-EEG, a novel EEG benchmark designed to advance research in visual neuroscience. CrossPT-EEG has a comprehensive and richly annotated dataset, comprising EEG recordings from 16 subjects, with a total of 22 sessions exposed to 4000 images from 80 different categories. We further establish benchmarks that encompass classification tasks on non-block-design EEG data and pre-training experiments spanning both time and participants. These benchmarks provide a systematic framework for evaluating the robustness and generalization of visual decoding models, and highlight the potential to advance cross-time, cross-subject learning in both cognitive neuroscience and brain–computer interface applications.
@@
Limitation. Firstly, while our dataset is more comprehensive than similar works, each participant’s data is still relatively limited. This necessitates the development of inter-subject models to overcome this limitation and enhance generalizability. Secondly, it is limited in representation, as participants were drawn from a specific sample. This results in an age distribution skewed towards teenager individuals and a racial composition predominantly White and Asian. Future work should aim to include a more diverse and extensive participant pool. Additionally, although our non-block-design paradigm and cross-time training strategy circumvent the temporal effects that often confound block-design experiments, we did not conduct an in-depth investigation of the temporal effects present in Stage 1. Future work could leverage our dataset to explore this direction more systematically. Lastly, our benchmarks did not incorporate many of the latest deep-learning methods. Instead, we employed a set of commonly used models primarily to demonstrate the effectiveness and to illustrate how our two-stage, cross-time dataset can be utilized. We believe that recent advancements in deep learning could greatly benefit from our comprehensive dataset, potentially leading to significant breakthroughs in visual neuroscience.
@@
Insight for ML. CrossPT-EEG provides a comprehensive resource for developing models in visual recognition tasks, enabling the development of sophisticated deep-learning models capable of capturing intricate patterns within EEG data. Future research could leverage the dataset to enhance domain adaptation and transfer learning techniques, facilitating effective inter-subject and cross-time task completion. By offering a diverse dataset of visual stimuli and supporting multi-level classification tasks, CrossPT-EEG could foster the creation of hierarchical models that mirror human cognitive processes and improve the generalization capabilities of machine learning algorithms. We hope this benchmark will enable the development of more sophisticated models and methodologies, driving forward EEG-based visual neuroscience research and offering deeper insights into the neural mechanisms underlying visual perception and processing.
@@
Insight for BCI. As hardware technology progresses, portable EEG devices are becoming increasingly feasible, offering new opportunities for real-time BCI applications. Researchers could develop robust BCI systems that accurately interpret user intent from EEG signals. The comprehensive size and diverse visual stimuli in CrossPT-EEG allow for the creation of adaptive BCI systems that learn and respond to long-term individual user patterns. This paves the way for personalized neurotechnology solutions, particularly enhancing human-computer interaction for individuals with disabilities. Furthermore, addressing privacy protection and ethical concerns will be crucial as BCI technology advances, ensuring user data is securely handled and individual rights are respected.
7 Ethics and Privacy
To protect participants’ privacy and physical health, our user study adheres to strict ethical guidelines for human research, with approval from the ethics committee111detailed information will be released after review. In accordance with ethical standards, we have taken several steps to protect participants’ privacy, including data anonymization and obtaining informed consent from all participants. Participants were fully briefed on the study’s aims, procedures, and potential implications. Moreover, the EEG recording procedure used is entirely non-invasive and involves no risk to participants.
8 Reproducibility statement
We are committed to full reproducibility of this work. The dataset in CrossPT-EEG will be publicly released after the review stage. All code, including data preprocessing, model training, and evaluation scripts, will be made available at github link. The architectural and implementation details of all models are documented in the Appendix A.3.
@@ A.1 Additional Information about Dataset
| CrossPT-EEG | 80 | 4000 | 16 | 87,850 | 20.85GB |
As shown in Listing 1, the CrossPT-EEG dataset storage format is provided after review. The dataset can be accessed through the cloud storage link available in our GitHub repository after review. Due to file size limitations on the cloud storage platform, we split the dataset of Stage 1 into two parts: “CrossPT-EEG_1.pth” and “CrossPT-EEG_2.pth”. Users can choose to use only one of the parts based on their specific needs or device limitations. Demographic information is also provided at the file level.